# On Finding Small Hyper-Gradients in Bilevel Optimization: Hardness Results and Improved Analysis

### v4

@@ 4.1 The Assumptions for Nonconvex-PL Bilevel Problems
Secondly, the stability of $Y_{\sigma}^{*}$ under perturbations of $\sigma$ by Lemma 4.1 implies the stability of $\varphi_{\sigma}$ and $\nabla\varphi_{\sigma}$ by invoking the result by Kwon et al. [29].
Lemma 4.3 (Kwon et al. [29]).
Recall that $\varphi(x)$ is the original hyper-objective in Equation 2, while $\varphi_{\sigma}(x)$ is the penalized hyper-objective in Equation 6. Under Assumption 4.1, $\nabla\varphi(x)$ exists and can be defined as the limit $\lim_{\sigma\rightarrow 0^{+}}\nabla\varphi_{\sigma}(x)$ Moreover, $\varphi_{\sigma}(x)$ is close to $\varphi(x)$. Formally, for any $0\leq\sigma\leq\min\{\rho_{g}/\rho_{f},\overline{\sigma}\}$, we have that
@@
[…]
@@
4.3 Near-Optimal Rates for the Deterministic Case
Interestingly, we can show a near-optima rate of $\tilde{\mathcal{O}}(\epsilon^{-2})$ by looking at the same algorithm from another viewpoint. Instead of viewing the algorithm as finding stationary points of an ${\mathcal{O}}(\sigma)$-close proxy $\varphi_{\sigma}(x)$, we view the algorithm as finding stationary points of the original hyper-objective $\varphi(x)$ with an ${\mathcal{O}}(\sigma)$-approximate gradient oracle $\nabla\varphi_{\sigma}(x)$. The benefit of this viewpoint is that the optimization process is carried out on $\varphi(x)$, which may have a smoother loss landscape than $\varphi_{\sigma}(x)$, as we can prove the following bound on the smoothness of $\varphi(x)$.
Lemma 4.4.
@@ 4.3 Near-Optimal Rates for the Deterministic Case
We sketch the proof as follows. First, we give the explicit form of $\nabla\varphi(x)$ in Lemma G.7. Next, we show that the PL condition ensures a strongly convex subspace near any minimum in Lemma G.6. Finally, We apply Lemma G.3 and G.5 to project the functions onto the strongly convex subspace, where $\varphi(x)$ has ${\mathcal{O}}(1)$-Lipschitz gradients to complete the proof.
@@
The above complexity matches the lower bound by Carmon et al. [6], except we require $f$ to be $\rho_{f}$-Hessian Lipschitz in Assumption 4.1d. However, as the dependency on $\rho_{f}$ is only logarithmic, we can still claim that our upper bound is near-optimal.
Remark 4.2.
Here we note some small differences between the assumptions in Theorem 4.1 and those of the constructed hard instance in Theorem 3.2.
@@
[…]
@@
4:   Set $K_{t}$ as Equation 8 based on the value of $\delta_{t}$
@@
11:   Calculate $\delta_{t+1}$ as Equation 9 based on $\delta_{t}$, $x_{t+1}$ and $x_{t}$
@@
Besides the near-optimal rate in the deterministic case, our novel viewpoint can also lead to better bounds on the stochastic case, where $f(x,y),g(x,y)$ are both the expectation of some stochastic components $F(x,y;\xi)$ and $G(x,y;\zeta)$, indexed by random vectors $\xi$ and $\zeta$:
@@
    Suppose the stochastic gradients have bounded variance. In other word, there exists some constants $M_{f},M_{g}>0$ such that
@@
[…]
Theorem 4.2.
@@
In the deterministic case $(M_{f}=0,M_{g}=0)$, this result recovers the near-optimal $\tilde{\mathcal{O}}(\epsilon^{-2})$ rate by Theorem 2. In the partially stochastic case ($M_{f}>0,M_{g}=0$), the $\tilde{\mathcal{O}}(\epsilon^{-4})$ complexity is also near-optimal [3]. In the fully stochastic case ($M_{f}>0,M_{g}>0$), our $\tilde{\mathcal{O}}(\epsilon^{-6})$ upper bound is also better than the $\tilde{\mathcal{O}}(\epsilon^{-7})$ upper bound by Kwon et al. [29]. But the bound still has a high dependency on $\ell,\kappa$, and $\epsilon$, and its tightness remains open.
@@ 5 Conclusions and Future Works
This paper investigates bilevel optimization without the typical lower-level strong convexity assumption. We have shown that finding a point with small a hyper-gradient is computationally hard for nonconvex-convex bilevel problems, but easy for nonconvex-PL bilevel problems where a simple first-order algorithm can achieve the near-optimal rate. In Appendix B, we discuss some future directions that have not been addressed by this work.
@@ Appendix A Related Works
Many existing works for non-asymptotic analysis for bilevel optimization assume the lower-level function is strongly convex. A natural method is to estimate $\nabla y^{*}(x)$, and plug into Equation 4 to get an estimator of $\nabla\varphi(x)$ to apply gradient descent. ITD (ITerative Differentiation) [19, 15, 48, 5] approximates $\nabla y^{*}(x)$ by ${\partial y^{K}(x)}/{\partial x}$, where $y^{K}(x)$ is $K$-steps of gradient descent. AID (Approximate Implicit Differentiation) [11, 17, 43, 16, 20, 24, 9, 1] explicitly solve $\nabla y^{*}(x)=\left[(\nabla_{yy}^{2}g)^{-1}\nabla_{yx}^{2}g\right](x,y^{*}(x))$ as a linear equation. However, both AID and ITD require Hessian-vector product oracles, and their convergence analysis is typically restricted to the case when the lower-level function is strongly convex. Arbel and Mairal [2] extended implicit differentiation to the parametric Morse-Bott function, a class of nonconvex functions with local PL properties. They also proposed an algorithm that combines AID and ITD and showed its limit points must be an equilibrium of the concept of BGS (Bilevel Game with Selection) that they introduced.
[…]
@@
[…]
GALET (Generalized ALternating mEthod for bilevel opTimization) by Xiao et al. [52] can also tackle nonconvex-PL bilevel optimization problems. It uses Hessian-vector-product oracles to solve the following gradient-based reformulation of Problem (1):
@@
Xiao et al. [52] showed GALET can converge to a KKT point of Problem (11) with a rate of $\tilde{\mathcal{O}}(\epsilon^{-2})$. In Appendix C, we show that this also implies the same convergence rate to the stationary point of $\varphi(x)$ under the assumptions of Kwon et al. [29].
Appendix B Future Works
Below, we discuss possible future directions.
Tightness of Results.
As we have discussed, we conjecture that the $\tilde{\mathcal{O}}(\epsilon^{-6})$ complexity in the fully stochastic setting is suboptimal and can be improved. Moreover, although we have established near-optimal complexity in both the deterministic setting and partially stochastic setting, our upper bounds still have some gap in logarithmic factors compared with the lower bounds. We hope the future works to close the gap. It is also interesting to see whether our dependency on $\kappa$ is (near)-optimal. A non-trivial combination of the existing lower bounds for minimizing nonconvex functions [6] in $x$ and PL functions [54] in $y$ may help this problem.
More Lower-Level Problems
It is also interesting to study bilevel problems with lower-level functions beyond the PL condition. One possible direction is to consider the Kurdyka-Łojasiewicz (KL) condition [13], which is more general than the PL condition and it holds for any semialgebraic functions. Recent works have shown non-asymptotic convergence for nonconvex-KL minimax optimization problems [30, 56], but the nonconvex-KL bilevel optimization problem remains challenging. Besides these conditions, we hope future works to exploit the problem structure that arises from the applications of bilevel optimization and identify other tractable conditions that can break the limit of our hardness result.
Appendix C Discussion on Xiao et al. [52]
The recently proposed GALET [52] has been shown to converge with the rate $\tilde{\mathcal{O}}(\epsilon^{-2})$ to the following definition of stationary points of Problem (11).
Definition C.1.
@@ Appendix B Future Works
Proposition C.1.
Under Assumption 4.1, if $x^{*}$ is an $\epsilon$-stationary point of $\varphi(x)$ in Equation 2, then it is also an ${\mathcal{O}}(\epsilon)$-stationary point as Definition C.1.
Proof.
Given $x^{*}$, take any $y^{*}\in Y^{*}(x)$, and let $w^{*}=-\left(\nabla_{yy}^{2}g(x^{*},y^{*})\right)^{\dagger}\nabla_{y}f(x^{*},y^{*})$. Then we have both ${\mathcal{R}}_{y}(x^{*},y^{*})=0$ and ${\mathcal{R}}_{w}(x^{*},y^{*},w^{*})=0$. By Lemma G.7, we know that ${\mathcal{R}}_{x}(x^{*},y^{*},w^{*})=\nabla\varphi(x^{*})$, therefore we also have $\|{\mathcal{R}}_{x}(x^{*},y^{*},w^{*})\|\leq\epsilon$.
@@
Proposition C.2.
Under Assumption 4.1 , if $\hat{x}$ is an $\epsilon$-stationary point as Definition C.1, then it is also an ${\mathcal{O}}(\epsilon)$-stationary point of $\varphi(x)$.
Proof.
Let $y^{*},w^{*}$ be the corresponding point of $x^{*}$ in Definition C.1. By Lemma G.2,
@@
${\mathcal{L}}(w)$ has the form of "strongly convex composed with linear". It is $s^{2}$-PL by Karimi et al. [26], Appendix B, where $s$ is the smaller singular value of $\nabla^{2}g(x^{*},\hat{y})$. By Lemma G.6, $s\geq\mu$.
Let $W^{*}=\arg\min_{w\in{\mathbb{R}}^{d_{y}}}{\mathcal{L}}(w)$. Then by Lemma G.1,
@@
Since $w^{*}$ is bounded by Lemma G.6, we know that
@@
Remark C.1.
In above analysis, the optimal $w^{*}$ of GALET [52] is the solution to the linear equation $\nabla_{yy}^{2}g(x,y^{*})w=\nabla_{y}g(x,y^{*})$. However, the Hessian matrix $\nabla_{yy}^{2}g(x,y)$ may be indefinite when the lower-level function is nonconvex. To overcome this issue, Xiao et al. [52] uses the square trick like Liu and Luo [36], which solves Equation 12 instead. One drawback of the square trick is the conditional number becomes ${\mathcal{O}}(\kappa^{2})$ and makes the inner loop slower. Since $\varphi(x)$ has ${\mathcal{O}}(\ell\kappa^{3})$-Lipschitz gradients by Lemma 4.4. According to our above analysis, the reasonable complexity of GALET should be $\tilde{\mathcal{O}}(\ell\kappa^{5}\epsilon^{-2})$ though the dependency on $\kappa$ is not explicitly given by Xiao et al. [52].
Remark C.2.
Xiao et al. [52] additionally assumes the smallest singular value of $\nabla_{yy}^{2}g(x,y)$ has a constant gap between zero when solving Equation 12. However, the complexity to solve Equation 12 only depends on the singular value of $\nabla_{yy}^{2}g(x,y^{*})$, which can be bound below by our Lemma G.6. Since $g(x,y)$ is PL in $y$, by Lemma G.8 gradient descent on $g$ can make ${\rm dist}(y^{k},Y^{*}(x))$ converges linearly. So one can let $y^{K}$ sufficiently close to some point in $Y^{*}(x)$ with logarithmic large $K$. Then by Weyl’s inequality the singular value of $\nabla^{2}_{yy}g(x,y^{K})$ is also bounded below. Therefore, such an additional assumption in Xiao et al. [52] seems to be redundant.
Appendix D Discussion on Kwon et al. [29]
Different from our Assumption 4.1a, Kwon et al. [29] uses the following proximal error bound (Prox-EB) assumption.
Assumption D.1.
@@
Proposition D.1.
Let $0\leq\sigma\leq\min\{L_{g}/L_{f},\overline{\sigma}\}$. Suppose both c and d in Assumption 4.1 hold. Let $\rho<1/(2L_{g})$. If Assumption 4.1a holds with constant $\mu$, then Assumption D.1 holds with $\mu^{\prime}=\mu/(1+2L_{g}\rho)$. Conversely, if Assumption D.1 holds with constant $\mu^{\prime}$, then Assumption 4.1a holds with $\mu=\left(\mu^{\prime}(1-2L_{g}\rho)\right)^{2}/(2L_{g})$.
Proof.
@@
Therefore, if $\mu$-PL condition holds, by Lemma G.1, we have
@@
Appendix E Proof of Theorem 3.1
See 3.1
Proof.

### v6

@@ 4.2 F2BA Can also be Applied to Nonconvex-PL Bilevel Problems
Secondly, the stability of $Y_{\sigma}^{*}$ under perturbations of $\sigma$ by Lemma 2 implies the stability of $\varphi_{\sigma}$ and $\nabla\varphi_{\sigma}$ by invoking the result by Kwon et al. (2024b).
Lemma 4 (Kwon et al. (2024b)).
Recall that $\varphi(x)$ is the original hyper-objective in Equation (2), while $\varphi_{\sigma}(x)$ is the penalized hyper-objective in Equation (6). Under Assumption 4.1, $\nabla\varphi(x)$ exists and can be defined as the limit $\lim_{\sigma\rightarrow 0^{+}}\nabla\varphi_{\sigma}(x)$ Moreover, $\varphi_{\sigma}(x)$ is close to $\varphi(x)$. Formally, for any $0\leq\sigma\leq\min\{\rho_{g}/\rho_{f},\overline{\sigma}\}$, we have that
@@
These two lemmas make it reasonable to apply the F2BA (Algorithm 2), which applies gradient descent on $\varphi_{\sigma}(x)$ according to Equation (7). Lemma 4 shows that to find an ${\mathcal{O}}(\epsilon)$-stationary point of $\varphi(x)$, it suffices to finds an ${\mathcal{O}}(\epsilon)$-stationary point of $\varphi_{\sigma}(x)$ for $\sigma={\mathcal{O}}(\epsilon)$. Note that gradient descent can find an $\epsilon$-stationary point of a nonconvex $L$-smooth function within ${\mathcal{O}}(L\epsilon^{-2})$ complexity and that $\varphi_{\sigma}(x)$ is ${\mathcal{O}}(\sigma^{-1})$-smooth in the worst-case. Such an analysis illustrates an $\tilde{\mathcal{O}}(\sigma^{-1}\epsilon^{-2})=\tilde{\mathcal{O}}(\epsilon^{-3})$ complexity as Kwon et al. (2024b).
Algorithm 2 F2BA $\left(x_{0},y_{0},\eta,\tau,\sigma,T,K\right)$
@@
4.3 Achieving the $\tilde{\mathcal{O}}(\epsilon^{-2})$ Rate for the Deterministic Case
Interestingly, we can show a rate of $\tilde{\mathcal{O}}(\epsilon^{-2})$ of Algorithm 2. Our improvement comes from a similar technique by Chen et al. (2023), which shows that $\varphi(x)$ is ${\mathcal{O}}(1)$-smooth for nonconvex-strongly-convex problems. The intuition is that we can restrict our analysis to the strongly convex subspace induced by the PL condition, and apply the result of Chen et al. (2023).
Lemma 5.
@@ 4.3 Achieving the $\tilde{\mathcal{O}}(\epsilon^{-2})$ Rate for the Deterministic Case
We sketch the proof as follows. First, we give the explicit form of $\nabla\varphi(x)$ in Lemma 7. Next, we show that the PL condition ensures a strongly convex subspace near any minimum in Lemma 6. Finally, We apply Lemma 3 and 5 to project the functions onto the strongly convex subspace, where $\varphi(x)$ has ${\mathcal{O}}(1)$-Lipschitz gradients to complete the proof.
@@
The above complexity matches gradient descent for single-level nonconvex minimization by Carmon et al. (2021), except we require $f$ to be $\rho_{f}$-Hessian Lipschitz in Assumption 4.1d. However, our dependency on $\rho_{f}$ is only logarithmic, rather than polynomial.
Remark 4.2
Here we note some small differences between the assumptions in Theorem 6 and those of the constructed hard instance in Theorem 2.
@@
4.4 The Extensions of F2BA to the Stochastic Case
Algorithm 3 F2BSA $\left(x_{0},y_{0},\delta_{0},\eta,\tau,\sigma,T,B\right)$
@@ 4.4 The Extensions of F2BA to the Stochastic Case
4:   Set $K_{t}$ as Equation (8) based on the value of $\delta_{t}$
@@
11:   Calculate $\delta_{t+1}$ as Equation (9) based on $\delta_{t}$, $x_{t+1}$ and $x_{t}$
@@
Our analysis can also lead to better bounds on the stochastic case, where $f(x,y),g(x,y)$ are both the expectation of some stochastic components $F(x,y;\xi)$ and $G(x,y;\zeta)$, indexed by random vectors $\xi$ and $\zeta$:
@@
    Suppose the stochastic gradients have bounded variance. In other words, there exist some constants $M_{f},M_{g}>0$ such that
@@
By replacing all the full-batch gradients in deterministic Algorithm 2, we get the stochastic counterpart Algorithm 3, namely F2BSA (Fully First-order Bilevel Stochastic Approximation). By additionally taking into account the error from stochastic gradients, we can extend Theorem 6 to also tackle the stochastic case, yielding the following theorem.
Theorem 7.
@@
In the deterministic case $(M_{f}=0,M_{g}=0)$, this result recovers the $\tilde{\mathcal{O}}(\epsilon^{-2})$ rate by Algorithm 2. In the partially stochastic case ($M_{f}>0,M_{g}=0$), the $\tilde{\mathcal{O}}(\epsilon^{-4})$ complexity also matches SGD (Arjevani et al., 2023). In the fully stochastic case ($M_{f}>0,M_{g}>0$), our $\tilde{\mathcal{O}}(\epsilon^{-6})$ upper bound is also better than the $\tilde{\mathcal{O}}(\epsilon^{-7})$ upper bound by Kwon et al. (2024b).
@@ 5 Conclusions and Future Works
This paper investigates bilevel optimization without the typical lower-level strong convexity assumption. We have shown that finding points with small hyper-gradients is computationally hard for nonconvex-convex bilevel problems, but easy for nonconvex-PL bilevel problems where simple first-order algorithms can achieve fast convergence rates. Finally, we discuss possible future directions.
Problem Complexity.
Assumption 4.1d suppose $f$ has $\rho_{f}$-Lipschitz Hessians, which comes from Lemma 4. Such an assumption is not required by nonconvex-strongly-convex problems (Chen et al., 2023). We wonder if it is possible to remove this assumption in our setting. Also, it is also interesting to study lower bounds for this problem (Kwon et al., 2024a).
More Lower-Level Problems
It is also interesting to study bilevel problems with lower-level functions beyond the PL condition. One possible direction is to consider the Kurdyka-Łojasiewicz (KL) condition (Fatkhullin et al., 2022), which is more general than the PL condition and it holds for any semialgebraic functions. Recent works have shown non-asymptotic convergence for nonconvex-KL minimax optimization problems (Li et al., 2022; Zheng et al., 2023), but the nonconvex-KL bilevel optimization problem remains challenging. Besides these conditions, we hope future works to exploit the problem structure that arises from the applications of bilevel optimization and identify other tractable conditions that can break the limit of our hardness result.
@@ Appendix A Related Works
Many existing works for non-asymptotic analysis for bilevel optimization assume the lower-level function is strongly convex. A natural method is to estimate $\nabla y^{*}(x)$, and plug into Equation (4) to get an estimator of $\nabla\varphi(x)$ to apply gradient descent. ITD (ITerative Differentiation) (Gould et al., 2016; Franceschi et al., 2017; Shaban et al., 2019; Bolte et al., 2021) approximates $\nabla y^{*}(x)$ by ${\partial y^{K}(x)}/{\partial x}$, where $y^{K}(x)$ is $K$-steps of gradient descent. AID (Approximate Implicit Differentiation) (Domke, 2012; Ghadimi and Wang, 2018; Pedregosa, 2016; Franceschi et al., 2018; Grazzi et al., 2020; Ji et al., 2021; Dagréou et al., 2022; Arbel and Mairal, 2021) explicitly solve $\nabla y^{*}(x)=\left[(\nabla_{yy}^{2}g)^{-1}\nabla_{yx}^{2}g\right](x,y^{*}(x))$ as a linear equation. However, both AID and ITD require Hessian-vector product oracles, and their convergence analysis is typically restricted to the case when the lower-level function is strongly convex. Arbel and Mairal (2022) extended implicit differentiation to the parametric Morse-Bott function, a class of nonconvex functions with local PL properties. They also proposed an algorithm that combines AID and ITD and showed its limit points must be an equilibrium of the concept of BGS (Bilevel Game with Selection) that they introduced. Yang et al. (2024) proposed a first-order algorithm FdeHBO by estimating Jacobian/Hessian vector-product in AID with gradient differences.
Recently, F2BA (Fully First-order Bilevel Approximation) has gained widespread attention. This method uses the value-function-based reformulation (Ye et al., 1997) of Problem (1):
@@
and applies gradient descent on the penalty function. Several studies have demonstrated two notable advantages of this method over AID and ITD: Firstly, it only requires gradient oracles which are much cheaper than Hessian-vector product oracles (Liu et al., 2022a). Secondly, it performs well empirically even in cases where the lower-level function is nonconvex (Liu et al., 2021a; Shen and Chen, 2023). By establishing the equivalence between the KKT (Karush–Kuhn–Tucker) point of Problem (10) and the stationary point of Problem (2), Kwon et al. (2023) showed that F2BA can find an $\epsilon$-stationary point of $\varphi(x)$ with $\tilde{\mathcal{O}}(\epsilon^{-3})$ first-order oracle calls when the lower-level function is strongly convex. Chen et al. (2023) further improved the rate to $\tilde{\mathcal{O}}(\epsilon^{-2})$ by a more careful landscape analysis of the penalty function. Kwon et al. (2024b); Yao et al. (2024) proposed proximal variants of F2BA by replacing the lower-level value function with its Moreau envelope to tackle lower-level constraints. Kwon et al. (2024b) showed a $\tilde{\mathcal{O}}(\epsilon^{-3})$ complexity when the penalty function satisfies the Prox-EB condition (which is equivalent to the PL condition in the unconstrained case as Proposition 1).
GALET (Generalized ALternating mEthod for bilevel opTimization) by Xiao et al. (2023) can also tackle nonconvex-PL bilevel optimization problems. It uses Hessian-vector-product oracles to solve the following gradient-based reformulation of Problem (1):
@@
Xiao et al. (2023) showed GALET can converge to a KKT point of Problem (11) with a rate of $\tilde{\mathcal{O}}(\epsilon^{-2})$. In Appendix B, we show that this also implies the same convergence rate to the stationary point of $\varphi(x)$ under the assumptions of Kwon et al. (2024b).
Huang (2023) also considers nonconvex-PL bilevel problems. Our work differs from (Huang, 2023) in the following aspects. First, Assumption 2 in (Huang, 2023) supposes that $\nabla^{2}g(x,y)$ is non-singular at the minimizer of g, while we do not require such an assumption. Secondly, our algorithm only requires first-order oracles, while (Huang, 2023) requires second-order oracles. Thirdly, our algorithm has only ${\mathcal{O}}(d)$ complexity at each step, where $d=\max\{d_{x},d_{y}\}$ is the dimension of the problem. In contrast, each iteration of the algorithm in (Huang, 2023) requires a ${\mathcal{O}}(d^{3})$ complexity for computing the SVD decomposition of the Hessian matrix.
Some works use a sequential approximation minimization strategy (Liu et al., 2021b, a, 2020) to tackle the discontinuous challenge for bilevel problems without lower-level strong convexity as illustrated by Example 3.1. These works generate a series of continuous functions $\{\varphi_{K}\}_{K=0}^{\infty}$ which converge to $\varphi$ when $K\rightarrow\infty$. Liu et al. (2020) proposed BDA by defining $\varphi_{K}(x)=f(x,{\rm AggrGD}_{f,g}^{K}(x))$ for nonconvex-convex problems, where ${\rm AggrGD}_{f,g}^{K}(x)$ denotes $K$-steps of gradient descent on the aggregated function $\alpha f(x,\,\cdot\,)+(1-\alpha)g(x,\,\cdot\,)$ for some $\alpha\in(0,1)$. Liu et al. (2021b) defined $\varphi_{K}(x)=\min_{z}\max_{k\leq K}\{f(x,{\rm GD}^{k}(x;z))\}$ for nonconvex-nonconvex problems, where ${\rm GD}^{k}(x;z))$ denotes $k$-steps of gradient descent on $g(x,\,\cdot\,)$ initialized with $z$. However, all these works require solving a series of sub-problems and often lack a non-asymptotic analysis.
Appendix B Discussion on Xiao et al. (2023)
The recently proposed GALET (Xiao et al., 2023) has been shown to converge with the rate $\tilde{\mathcal{O}}(\epsilon^{-2})$ to the following definition of stationary points of Problem (11).
Definition B.1
@@
Proposition 1.
Under Assumption 4.1, if $x^{*}$ is an $\epsilon$-stationary point of $\varphi(x)$ in Equation (2), then it is also an ${\mathcal{O}}(\epsilon)$-stationary point as Definition B.1.
Proof B.1.
Given $x^{*}$, take any $y^{*}\in Y^{*}(x)$, and let $w^{*}=-\left(\nabla_{yy}^{2}g(x^{*},y^{*})\right)^{\dagger}\nabla_{y}f(x^{*},y^{*})$. Then we have both ${\mathcal{R}}_{y}(x^{*},y^{*})=0$ and ${\mathcal{R}}_{w}(x^{*},y^{*},w^{*})=0$. By Lemma 7, we know that ${\mathcal{R}}_{x}(x^{*},y^{*},w^{*})=\nabla\varphi(x^{*})$, therefore we also have $\|{\mathcal{R}}_{x}(x^{*},y^{*},w^{*})\|\leq\epsilon$.
@@
Proposition 2.
Under Assumption 4.1 , if $\hat{x}$ is an $\epsilon$-stationary point as Definition B.1, then it is also an ${\mathcal{O}}(\epsilon)$-stationary point of $\varphi(x)$.
Proof B.2.
Let $y^{*},w^{*}$ be the corresponding point of $x^{*}$ in Definition B.1. By Lemma 2,
@@
${\mathcal{L}}(w)$ has the form of ”strongly convex composed with linear”. It is $s^{2}$-PL by Karimi et al. (2016), Appendix B, where $s$ is the smaller singular value of $\nabla^{2}g(x^{*},\hat{y})$. By Lemma 6, $s\geq\mu$.
Let $W^{*}=\arg\min_{w\in{\mathbb{R}}^{d_{y}}}{\mathcal{L}}(w)$. Then by Lemma 1,
@@
Since $w^{*}$ is bounded by Lemma 6, we know that
@@
Remark B.1
In above analysis, the optimal $w^{*}$ of GALET (Xiao et al., 2023) is the solution to the linear equation $\nabla_{yy}^{2}g(x,y^{*})w=\nabla_{y}g(x,y^{*})$. However, the Hessian matrix $\nabla_{yy}^{2}g(x,y)$ may be indefinite when the lower-level function is nonconvex. To overcome this issue, Xiao et al. (2023) uses the square trick like Liu and Luo (2022), which solves Equation (12) instead. One drawback of the square trick is the conditional number becomes ${\mathcal{O}}(\kappa^{2})$ and makes the inner loop slower. Since $\varphi(x)$ has ${\mathcal{O}}(\ell\kappa^{3})$-Lipschitz gradients by Lemma 5. According to our above analysis, the reasonable complexity of GALET should be $\tilde{\mathcal{O}}(\ell\kappa^{5}\epsilon^{-2})$ though the dependency on $\kappa$ is not explicitly given by Xiao et al. (2023).
Remark B.2
Xiao et al. (2023) additionally assumes the smallest singular value of $\nabla_{yy}^{2}g(x,y)$ has a constant gap between zero when solving Equation (12). However, the complexity to solve Equation (12) only depends on the singular value of $\nabla_{yy}^{2}g(x,y^{*})$, which can be bound below by our Lemma 6. Since $g(x,y)$ is PL in $y$, by Lemma 8 gradient descent on $g$ can make ${\rm dist}(y^{k},Y^{*}(x))$ converges linearly. So one can let $y^{K}$ sufficiently close to some point in $Y^{*}(x)$ with logarithmic large $K$. Then by Weyl’s inequality the singular value of $\nabla^{2}_{yy}g(x,y^{K})$ is also bounded below. Therefore, such an additional assumption in Xiao et al. (2023) seems to be redundant.
Appendix C Discussion on Kwon et al. (2024b)
Different from our Assumption 4.1a, Kwon et al. (2024b) uses the following proximal error bound (Prox-EB) assumption.
Assumption C.1
@@
Proposition 1.
Let $0\leq\sigma\leq\min\{L_{g}/L_{f},\overline{\sigma}\}$. Suppose both c and d in Assumption 4.1 hold. Let $\rho<1/(2L_{g})$. If Assumption 4.1a holds with constant $\mu$, then Assumption C.1 holds with $\mu^{\prime}=\mu/(1+2L_{g}\rho)$. Conversely, if Assumption C.1 holds with constant $\mu^{\prime}$, then Assumption 4.1a holds with $\mu=\left(\mu^{\prime}(1-2L_{g}\rho)\right)^{2}/(2L_{g})$.
Proof C.1.
@@
Therefore, if $\mu$-PL condition holds, by Lemma 1, we have
@@
Appendix D Proof of Theorem 1
See 1
Proof D.1.